OpenStudy (anonymous):

Using Pascal's Triangle, expand (x+3y)^3 and (m+3n)^3. Show work.

OpenStudy (anonymous):

(a + b)^3 = a^3 + 3(a^2)b + 3a(b^2) + b^3. For the first problem a = x and b = 3y. Similar for the second with a change of variables.

OpenStudy (anonymous):

So far I have x^3+9x^2y but I'm not sure what else to do.

OpenStudy (anonymous):

Hold on and I'll upload a picture.

OpenStudy (anonymous):

This what I have so far.

OpenStudy (anonymous):

You're on the right track, but you need to be including the "3" when you are putting an exponent on "y". It is NOT understood to extend to the 3. 3y^2 is not the same as (3y)^2. That will mess you up eventually and is wrong notation. Use the parentheses.

OpenStudy (anonymous):

So, do I square 3 which gives me 9 then multiply it by 3 to get 27? Then I would have x^3+9x^2y+27xy^2 so far right?

OpenStudy (anonymous):

Absolutely correct! You are understanding this well. It's just your notation that was a little off before. I think you are solidly on track now.

OpenStudy (anonymous):

Is the final answer x^3+9x^2y+27xy^2+27y^3?

OpenStudy (anonymous):

You're right there! 100% ! I've helped you on a different question and you do very well. Good work!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You are quite welcome!